OpenStudy (anonymous):

In triangle ABC, a=6, b=9 and sin A=2/3. Find sin B

OpenStudy (anonymous):

the "Law of sines" tells you, that \[{\sin \alpha \over a} = {\sin \beta \over b} = {\sin \gamma \over c} \]

OpenStudy (anonymous):

ok